# Detection of New Galaxy Candidates at 11 in the JADES Field Using JWST NIRCam

### v2

@@ 5 Conclusion
In this paper, we present the outcomes of our investigation aimed at identifying z>11 galaxy candidates in the JWST NIRCam observations of JADES GOODS-S and GOODS-N fields. We undertake an extensive data reduction process that involves several custom procedures. We use two distinct codes, EAZY and BAGPIPES, for photometric redshift determination and investigating the properties of these galaxies. We detect seven z>11 candidate galaxies distributed as follows: three candidates between 11 < z < 12, and four candidates between 12 < z < 15. All candidates are robust high-redshift detections with 70-100% of their probability within , meeting the classification ‘high-quality’ used in a recent JWST UNCOVER high-redshift survey (Atek et al., 2023b). The detection of these young galaxies spans a diverse range of magnitudes, with F277W AB Kron magnitude values between 28.5 and 30.4, consistent with previous JADES surveys, which achieved detections as faint as AB mag 31.0 for the young galaxies at redshifts greater than 10 (Hainline et al., 2024).
To assess the physical properties of the galaxies, we perform SED fitting to the photometric data of these galaxies using BAGPIPES. We obtain measurements for stellar mass, SFRs averaged over the last 100 Myr, specific star sSFR, dust content, and mass-weighted age, which are documented in Table 1. We find the intrinsic stellar masses of our galaxy candidates to vary within the range of to . Their SFRs span between 0.9 and 5.1 , while their sSFRs, quantified as , range from 0.95 to 1.46. The mass-weighted ages of these galaxies range between 80 to 240 Myr, with their dust content varying between = 0.02 and 0.45. We present the distribution of galaxy candidates on the stellar mass versus star formation rate plane and compare our findings, which closely align with those of previous surveys (Nakajima et al., 2023; Morales et al., 2024; Curti et al., 2024).
@@ Acknowledgements
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. We thank Dr. Daniel Eisenstein for the helpful discussion on JWST/JADES data. We acknowledge support by NASA ADAP Grant 80NSSC21K0637.

### v3

@@ 5 Conclusion
In this paper, we present the outcomes of our investigation aimed at identifying galaxy candidates in the JWST NIRCam observations of JADES GOODS-S and GOODS-N fields. We undertake an extensive data reduction process that involves several custom procedures. We use two distinct codes, EAZY and BAGPIPES, for photometric redshift determination and investigating the properties of these galaxies. We detect six candidate galaxies distributed as follows: three candidates between , and four candidates between . All candidates are robust high-redshift detections with 70-100% of their probability within , meeting the classification ‘high-quality’ used in a recent JWST UNCOVER high-redshift survey (Atek et al., 2023a). The detection of these young galaxies spans a diverse range of magnitudes, with F277W AB Kron magnitude values between 28.5 and 30.4, consistent with previous JADES surveys, which achieved detections as faint as AB mag 31.0 for the young galaxies at redshifts greater than 10 (Hainline et al., 2024).
To assess the physical properties of the galaxies, we perform SED fitting to the photometric data of these galaxies using BAGPIPES. We obtain measurements for stellar mass, SFRs averaged over the last 100 Myr and 10 Myr, dust content, and mass-weighted age, which are listed in Table 1. We find the intrinsic stellar masses of our galaxy candidates to vary within the range of to . Their SFRs span between 0.9 and 5.1 , while their sSFRs, quantified as , range from 0.95 to 1.46. The mass-weighted ages of these galaxies range between 80 to 240 Myr, with their dust content varying between = 0.02 and 0.45. We present the distribution of galaxy candidates on the stellar mass versus star formation rate plane and compare our findings, which closely align with those of previous surveys (Nakajima et al., 2023; Morales et al., 2024; Curti et al., 2024).
@@ Acknowledgements
We thank the anonymous referee for their helpful comments and suggestions. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. We thank Dr. Daniel Eisenstein for the helpful discussion on JWST/JADES data. We acknowledge support by NASA Chandra grant GO5-26101X 16619404 and NASA ADAP Grant 80NSSC21K0637.